OpenStudy (anonymous):

In a certain county, the number of Charter Schools is ten less than twice the number of Alternative Schools. We know that there are 46 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (anonymous):

Can you develop and equation to reprsent this situation?

OpenStudy (anonymous):

*an equation

OpenStudy (anonymous):

(And I'm asking as in can you have a try at doing it yourself before I/somone else jumps in to solve it for you;) )

OpenStudy (anonymous):

i tried .. 46 = 10 - 2x but i dont get how to get X by itself to find the value

OpenStudy (anonymous):

28 alternative schools.

OpenStudy (anonymous):

2x-10=y x=alternative schools, y=charter schools 2x-10=46 2x=56 x=28

OpenStudy (anonymous):

Close, but it's the other way around. Let A = alternative, C = charter. C = 2A - 10 C = 46 2A - 10 = 46 Now try grouping like terms. Move all the terms containing the variable to one side, and all the other terms to the other side. Remember that what you do on one side must be done on the other; say you subtract 50 from one side, the other side must gain 50 (it can't simply disappear).

OpenStudy (anonymous):

I'm trying to explain the concept here babyzoe. Working out algebra isn't hard for those who know, but learning it can be difficult.

OpenStudy (anonymous):

ooooo im sorry but i did show her hw to get it

OpenStudy (anonymous):

if only i had teachers like you two i would get it.. thanks bunches i can do it now .. yay me!

OpenStudy (anonymous):

your welcome